OpenStudy (anonymous):

Last One , Please Help Me ? Solve, using the substitution method. y + 2x = 7 14 – 4x = 2y A. The solution is (1, 5) B. The solution is (21, 0) C. There are an infinite number of solutions. D. There is no solution.

OpenStudy (anonymous):

Ok then lets use the first equation. Leave the 'y' alone, like this: \[y = 7 -2x\] and plug this 'y' in the second equation. This method is called substitution. \[14 - 2x = 2(7 - 2x)\] \[14 - 2x = 14 - 4x\] Fourteens will simplify itselves, so you get : \[2 = 4\] So there is no solution.